# Quantum Complexity in Constrained Many-Body Models: Scars, Fragmentation, and Chaos

### v1

@@ IV.2.2 Level spacing ratio
IV.2.3 Spectral Form Factor
@@
Our analysis also shows the advantage of using the Spectral Form Factor (SFF) compared to level statistics. As noted earlier, the additional spin-flip symmetry in the model $H^{\mathcal{N}(2)}$ requires an extra step of symmetry resolution. Without this step, the level spacing distribution deviates strongly from the Wigner–Dyson form and instead resembles a Poisson distribution, as seen in Fig. 12(a) in Appendix C. In contrast, the SFF still displays the characteristic ramp behavior even without resolving this last symmetry. See Fig. 12(b). This suggests that, for quantum many-body models where symmetry resolution plays a crucial role in identifying genuine chaotic behavior, the SFF can provide a more reliable diagnostic.
IV.3 Stabilizer Rényi Entropy
The final aspect of quantum complexity discussed in our work is nonstabilizerness. A useful way to quantify nonstabilizerness is by measuring how far a quantum state is from being a stabilizer state, which is referred to as the Stabilizer Rényi Entropy (SRE). For a pure state $|\Psi\rangle$ of $L$ qubits, the SRE is defined as [31]:
In our study, we examine a maximum system size of $L=12$ and plot the SRE behavior for all the considered models in Fig. 6. Similar to the behavior of entanglement as illustrated in Figs. 3 and 4, the SRE value obtained for the symmetry resolved fragmented sectors of $H^{\mathcal{N}(1)}$ (a) and $H^{\mathcal{N}(2)}$ (b) shows a more spread behavior than the Quantum Game of Life (f), which shows smooth behavior and remains nearly flat with variations in energy. This flatness indicates that the eigenstates in the middle of the spectrum exhibit some variation in terms of entanglement while sharing the same SRE content. This behavior resembles the scenario where quantum states remain maximally localized in the Fock basis, resulting in flat Shannon information entropy as discussed in [58]. In case of $H^{\mathcal{N}(1)}$ and $H^{\mathcal{N}(2)}$, we separately plot the SRE behavior for a few of the largest subspaces to compare the SRE generating power of the dynamically disconnected Krylov sectors. Figure 6(d) reveals that for $H^{\mathcal{N}(1)}$, the highest and the second-highest dimensional subspaces generate almost the same SRE value. This is different from the entanglement behavior we observed in Fig. 3(c). Whereas, for $H^{\mathcal{N}(2)}$, Fig. 6(d) shows many of the eigenstates of the second-highest subspace from the middle of the spectrum show higher SRE values than those obtained for the highest-dimensional sector.
Figure 7: Behavior of return probability and overlap with energy eigenstates of an initial state $|K\rangle^{\otimes 2}$ as defined in Eq. (11) when quenched using Hamiltonian $H^{\mathcal{N}(0)}$. (a) shows a clear revival of the state signaturing existence of quantum scars that can also be validated further by taking the overlap of $|K=6\rangle^{\otimes^{2}}$ with all the eigenstates of the model as shown in (b). Here, we can clearly see that four eigenstates exist, sharing a relatively higher value of the overlap (marked using red circles), which characterizes the typical scar states. This behavior is almost the same as obtained in Fig. 1(a) of Ref. [41] for the conventional PPXPP model with Rydberg constraint. Both the plots are obtained for $L=12$.
@@ V Quantum scars in fragmented subspaces
Appearance of kinetically constrained models that consist of a set of atypical states that do not thermalize is not a new phenomena. In literature, the existence of such models have been well studied [13, 11, 12, 14] and recently, some prescriptions have also been suggested to find whether quantum many-body scar exists [59, 60]. However, for a given many-body model, finding scar is a nontrivial task. In our work, we start from the case where it is already known that scar exist and we focused on their robustness against nontrivial perturbations.
We build on the $H^{\mathcal{N}(0)}$ model or unconstrained PPXPP model. As Fig. 7 shows, the time evolution of $\langle\Psi(t)|\Psi(0)\rangle$ of the initial state $|\Psi(0)\rangle=|K=6\rangle^{\otimes 2}$ with
@@
which corresponds to a uniform superposition with a single excitation $|1\rangle$ across the $K/2$ sites of the second half of the unit cell. The figure shows strong revivals, signifying the existence of quantum scar states that we also confirm by taking the overlap of $|K=6\rangle^{\otimes^{2}}$ with the eigenstates of the model as depicted in Fig. 7(b). This results is very similar to that obtained for the conventional PPXPP model with Rydberg constraint as already reported in Fig. 1(a) of the Ref.[41].
Figure 8: Illustration of the existence of quantum scars in fragmented subspaces of the model $H^{\text{Pert}}_{\text{PPXPP}}$. Similar to the marker of quantum scars depicted in Fig. 7, we notice that (a) and (b) show similar behavior. In (c), we explicitly present the graphical representation of the fragmented subspaces, akin to what is shown in Fig. 1. Additionally, the spectral distribution is illustrated in (d), confirming the robustness of the chaotic nature of this model against perturbations. Here, for (a) and (b), we have considered the system size $L=12$. Whereas. the schematic in (c) is obtained for $L=10$. For (d), we considered $L=22$.
@@
Table 3: Relevant Hilbert space dimensions and sector counts for different system sizes $L$ obtained for $H_{\text{PPXPP}}^{\text{Pert}}=H^{\mathcal{N}(0)}+\delta\cdot H^{\mathcal{N}(1)}$. Here, for a system with size $L$, we denote the number of Krylov subspaces by $\mathcal{N}^{01}$, dimension of the largest Krylov subspace by $d_{L}^{01}$ and the dimension of the symmetry-resolved Hilbert space by $\mathcal{D}_{L}^{01}$.
We next study how far this scar behavior persists if we add some perturbation to the model. In particular, we added $H^{\mathcal{N}(1)}$ with the model as perturbation and studied similar things for $H^{\text{Pert}}_{\text{PPXPP}}=H^{\mathcal{N}(0)}+\delta\cdot H^{\mathcal{N}(1)}$. Interestingly, we note that such behavior persists for the value of $\delta\approx 0.09$, as can be seen in Fig. 8. The result has significant importance as we report such a model that has both non-trivial Hilbert space fragmentation (see Table 3 for the scaling of number of Krylov spaces ($\mathcal{N}_{L}$) and the dimension of the highest one ($d_{L}$) with system size $L$) as well as scar-like feature that is not present together in any other models that we have considered in our work. In particular, QGL does not show Hilbert space fragmentation and, as the entanglement entropy plot shows no atypical behavior, we conjecture the nonexistence of quantum many-body scar. However, as mentioned in Sec . III.3, although $H^{\mathcal{N}(0)}$ can be regarded as another model where quantum scars and fragmentation coexist, the number of non-trivial fragmented subspaces in this case is very limited, and under the Rydberg constraint they completely vanish. For this reason, we attribute the perturbed PPXPP model as a more suitable candidate to exhibit both phenomena simultaneously. It is also not straightforward to recover any existing scar-like feature for $H^{\mathcal{N}(1)}$ (or $H^{\mathcal{N}(3)}$) and $H^{\mathcal{N}(2)}$. We wish to dig deeper about this in future work for a formal proof of the existence or nonexistence of quantum scars in these models.
VI Quantum complexity study of $H_{\mathrm{Tot}}$
@@
Figure 9: Quantum complexity for the Hamiltonian $H_{\mathrm{Tot}}$, as defined in Eq. (1). (a) The behavior of entanglement for $L=18$ we obtain here remains notably distinct from that obtained for other KCMs discussed previously. Notably, in this case, in addition to the primary arch-like behavior, the presence of low-entangled states creates a sub-arch structure or an entanglement band. Similar patterns are observed in the plots shown in (b) and (c), where we plot the behavior of the Inverse Participation Ratio (IPR) for $L=18$ and the Stabilizer Rényi Entropy, $\mathcal{M}$ for $L=12$, respectively. In case of (b), the y-axis is rescaled by a factor $10^{-3}$. Finally, (d) illustrates the chaotic behavior of the same model for $L=20$.
@@
We devote this section to discuss the quantum complexity properties of the Hamiltonian $H_{\mathrm{Tot}}$, separately. Interestingly, in this case, we observe that although the model does not exhibit any fragmentation behavior, its quantum complexity behavior shows a distinct feature from all the other models we have discussed so far. The first distinct feature we highlight here is the EE behavior with energy as shown in Fig. 9(a), displaying sub-arch-like splitting of the primary arch shape, resulting in an entanglement band-like structure. This behavior we find counterintuitive, and it does not translate immediately from all other observed behavior. To substantiate our claim, in Fig. 9(b), we also plot the Inverse Participation Ratio (IPR), defined as
@@
with $\{|i\rangle\}$ being the computational basis. The figure shows that IPR exhibits the same behavior as EE. The same imprint can be found in SRE as depicted in Fig. 9(c). Hence, although the system does not exhibit conventional fragmentation, it indicates that EE and other measures are not smooth functions of energy. Even in the bulk, there exist eigenstates that cause dips in the entanglement and create this band-like structure. This certainly opens up the possibility of investigating such a behavior in more detail and exploring the connection with the existence of quantum scars in these models, which we plan to pursue in our future work. The symmetry-resolved sectors of the model show chaotic nature, as displayed in Fig. 9(d), which remains consistent with the behavior obtained in the previous results.
@@ VII Discussion
In this work, we have investigated quantum complexity in kinetically constrained models, examining both the challenges of state preparation and the growth of complexity under dynamical evolution. The family includes the celebrated quantum Game of Life and the unconstrained PPXPP model along with their variations. We have shown how entanglement, nonstabilizerness, and quantum chaos interplay with each other in these models. In particular, we have unveiled the chaotic behavior through systematic symmetry resolutions of the models. However, there are cases where mere symmetry resolution has not revealed the chaotic behavior, as the Hilbert space exhibits dynamically disconnected sectors, which has resulted in Hilbert space fragmentation. Along with that, for some of the models, the fragmented subspaces have further hosted quantum many-body scars. This has given us a scope to extract non-trivial physics that exhibits quantum chaos, quantum many-body scars, as well as Hilbert space fragmentation out of a single family of quantum many-body models.
We have further characterized the fragmented subspaces by their resource generation capabilities. In particular, we have studied how the dimension of the Krylov subspaces correlates with the maximum entanglement entropy generation as well as their classical simulability aspects as quantified by Stabilizer Rényi entropy. Our analysis has revealed that although the resource generation capabilities of the smaller Krylov subspaces are low, the highest entanglement may not always come from the largest Krylov subspace. Hence, our analysis has shed light on drastically different quantum properties of structurally similar kinetically constrained models that can be instrumental for further fundamental as well as application-based studies. For example, in our work on QGL and related models, we have not found evidence of quantum many-body scars, which we plan to pursue more systematically in future work. Along with that, whether the quantum many-body scar or the fragmented phase can be applied for practical applications such as quantum sensing is another direction we also wish to explore. Another aim will be to characterize the fragment subspaces in terms of their Krylov Complexity behavior [61, 62]. Finally, the entanglement localization feature imprinted in all the quantum properties we obtained for the model $H_{\mathrm{Tot}}$ is another aspect that we will study in detail in our future work.

### v2

@@ IV.3.2 Level spacing ratio
IV.3.3 Spectral Form Factor
@@
Our analysis also shows the advantage of using the Spectral Form Factor (SFF) compared to level statistics. As noted earlier, the additional spin-flip symmetry in the model $H^{\mathcal{N}(2)}$ requires an extra step of symmetry resolution. Without this step, the level spacing distribution deviates strongly from the Wigner–Dyson form and instead resembles a Poisson distribution, as seen in Fig. 18(a) in Appendix D. In contrast, the SFF still displays the characteristic ramp behavior even without resolving this last symmetry. See Fig. 18(b). This suggests that, for quantum many-body models where symmetry resolution plays a crucial role in identifying genuine chaotic behavior, the SFF can provide a more reliable diagnostic.
Figure 10: Behavior of return probability and overlap with energy eigenstates of an initial state $|K=6\rangle^{\otimes 2}$ as defined in Eq. (11) when quenched using Hamiltonian $H^{\mathcal{N}(0)}$. (a) shows a clear revival of the state signaturing existence of quantum scars that can also be validated further by taking the overlap of $|K=6\rangle^{\otimes 2}$ with all the eigenstates of the model as shown in (b). Here, we can clearly see that four eigenstates exist, sharing a relatively higher value of the overlap (marked using red circles), which characterizes the typical scar states. This behavior is exactly the same as obtained in Fig. 1(a) of Ref. [47] for the conventional PPXPP model with Rydberg constraint as the largest Krylov space in both the cases coincides. Both the plots here are obtained for $L=12$.
Figure 11: Illustration of the existence of quantum scars in fragmented subspaces of the model $H^{\text{Pert}}_{\text{PPXPP}}(k)$. Similar to the marker of quantum scars depicted in Fig. 10, we notice that (a) and (b) show similar behavior for $H^{\text{Pert}}_{\text{PPXPP}}(k=1)$ with $\delta=0.09$, and (c)-(d) depicts the same for $H^{\text{Pert}}_{\text{PPXPP}}(k=2)$ with $\delta=0.50$. Here we considered $L=12$.
@@ V Quantum scars in fragmented subspaces
Appearance of kinetically constrained models that consist of a set of atypical states that do not thermalize is not a new phenomena. In literature, the existence of such models have been well studied in PXP and related family of models [14, 12, 13, 15, 62, 63] and recently, some prescriptions have also been suggested to find whether quantum many-body scar exists [64, 65]. However, for a given many-body model, finding scar is a nontrivial task.
In our work, the entanglement behavior observed for $H^{\mathcal{N}(1)}$, $H^{\mathcal{N}(2)}$, $H^{\text{QGL}}$ and does not provide evidence for the presence of quantum many-body scars. Additional support for this conclusion comes from the absence of revival dynamics when standard initial states such as $|\mathbb{Z}_{2}\rangle$ and $|\mathbb{Z}_{4}\rangle$, commonly used in previous studies, are employed. However, we demonstrate that when these models are combined with $H^{\mathcal{N}(0)}$, signatures of persistent nonthermal dynamics survive even in the strong perturbation limit. Before presenting these results, we remind the reader that the largest Krylov subspace of $H^{\mathcal{N}(0)}$ coincides with the conventional PPXPP model with Rydberg constraints for which the characteristic features of quantum scar states are already known. For completeness, we first reproduce this behavior by examining the time evolution of the overlap $\langle\Psi(t)|\Psi(0)\rangle$ for an appropriate initial state evolved under $H^{\mathcal{N}(0)}$, as reported in Fig. 1(a) of Ref. [47]. Here, $|\Psi(0)\rangle=|K=6\rangle^{\otimes 2}$ with
@@
which corresponds to a uniform superposition with a single excitation $|1\rangle$ across the $K/2$ sites of the second half of the unit cell. Figure 10 shows strong revivals in the behavior of $\langle\Psi(t)|\Psi(0)\rangle$, signifying the existence of quantum scar states that we also confirm by taking the overlap of $|K=6\rangle^{\otimes 2}$ with the eigenstates of the model.
Figure 12: Intensity plot of the squared overlap between the exact eigenstates $|E\rangle$ of the perturbed Hamiltonian $H^{\text{Pert}}_{\text{PPXPP}}(k)$ and the exact scar state $|E_{s}\rangle$ of $H^{\mathcal{N}(0)}$ for (a) $k=1$ and (b) $k=2$, respectively. The results are consistent with those shown in Fig. 11. In both cases, beyond a critical value of the detuning parameter $\delta$ (e.g., $\delta\lesssim 0.09$ for $k=1$ and $\delta\lesssim 0.50$ for $k=2$), the scar disappears, as the probe state develops significant overlap with other eigenstates (away from the bulk) as well.
We next study the robustness of the scar behavior under perturbations. In particular, we perturb the model by adding $H^{\mathcal{N}(k)}$ and consider the Hamiltonian
$H^{\text{Pert}}_{\text{PPXPP}}(k)=H^{\mathcal{N}(0)}+\delta\cdot H^{\mathcal{N}(k)}.$ (12)
We find that the scar signatures persist up to finite perturbation strengths, namely $\delta\approx 0.09$ for $k=1$ and $\delta\approx 0.50$ for $k=2$, as shown in Fig. 11. The robustness of the scar is illustrated more explicitly in Fig. 12, where in both cases in the respective $\delta$ range (e.g., $\delta\lesssim 0.09$ for $k=1$ and $\delta\lesssim 0.50$ for $k=2$), a single eigenstate of the perturbed Hamiltonian can be continuously traced back to the scar state of the unperturbed model. This can be realized when we examine the squared overlap between the exact eigenstates of the perturbed Hamiltonian $H^{\text{Pert}}_{\text{PPXPP}}(k)$ ($|E\rangle$) and the exact scar state of $H^{\mathcal{N}(0)}$ ($|E_{s}\rangle$).
@@
Table 3: Relevant Hilbert space dimensions and Krylov sector counts for the perturbed Hamiltonian $H^{\text{Pert}}_{\text{PPXPP}}(k)=H^{\mathcal{N}(0)}+\delta\,H^{\mathcal{N}(k)}$. For a system of size $L$, $\mathcal{N}^{0k}_{L}$ denotes the number of Krylov subspaces, $d^{0k}_{L}$ the dimension of the largest Krylov subspace, and $\mathcal{D}^{0k}_{L}$ the symmetry-resolved Hilbert space dimension.
Hamiltonian Fragmentation Scarring
$H^{\mathcal{N}(0)}$ Trivial HSF Strong scars
$H^{\mathcal{N}(0)}+\delta\cdot H^{\mathcal{N}(1)}$ Strong HSF Weak scars
($\delta\lesssim 0.09$)
$H^{\mathcal{N}(0)}+\delta\cdot H^{\mathcal{N}(2)}$ Weak HSF Strong scars
($\delta\lesssim 0.50$)
Table 4: Summary of Hilbert-space fragmentation (HSF) and quantum many-body scarring behavior for the unperturbed and perturbed $H^{\mathcal{N}(0)}$ Hamiltonians. The table highlights the coexistence of fragmentation and scarring, as well as their distinct stability regimes under controlled perturbations. Here, weak scars denote scar signatures that vanish under arbitrarily small perturbations $\delta\gtrsim 0.09$ as can be seen for $H^{\mathcal{N}(1)}$, whereas for $H^{\mathcal{N}(2)}$ the scarring behavior remains robust and survives up to perturbation strengths of approximately $\delta\lesssim 0.5$. Similarly, as mentioned in the main text, here trivial HSF refers to the case where, although the condition $d_{L}/\mathcal{D}_{L}\rightarrow 0$ as $L\gg 1$ is formally satisfied, most of the Krylov subspaces consist of a single basis state that is trivially annihilated by the Hamiltonian due to the kinetic constraint.
These observations demonstrate that $H^{\text{Pert}}_{\text{PPXPP}}$ provides an explicit example of a system that simultaneously hosts Hilbert-space fragmentation and quantum many-body scars. Notably, as we observed from Fig. 12, the scarred eigenstates can be continuously traced back to the unperturbed limit, for some values of $\delta$, implying that, in terms of quantum scars, the perturbations do not qualitatively change the physics from the unperturbed case. However, the fragmentation behavior of $H^{\text{Pert}}_{\text{PPXPP}}$ remains qualitatively distinct from that of $H^{\mathcal{N}(0)}$, $H^{\mathcal{N}(1)}$, or $H^{\mathcal{N}(2)}$ as can be seen from Fig. 16 presented in Appendix B.
Table 4 and explicitly show the relevant Hilbert space dimensions and Krylov sector counts for the perturbed PPXPP model and helps us in categorizing these models in three cases: (i) $H^{\mathcal{N}(0)}$ exhibits strong scarring without nontrivial fragmentation (the Krylov subspaces are dominated by sectors consisting of a single basis state, which is annihilated by the Hamiltonian due to the kinetic constraint). (ii) $H^{\text{Pert}}_{\text{PPXPP}}(k=1)$ shows strong fragmentation with weak but stable scarring, which survives up to $\delta\lesssim 0.09$. (iii) $H^{\text{Pert}}_{\text{PPXPP}}(k=2)$ displays relatively weaker fragmentation than $H^{\text{Pert}}_{\text{PPXPP}}(k=1)$ together with comparatively more robust scarring, persisting up to $\delta\lesssim 0.50$. This establishes a regime in which three distinct features: scarring, fragmentation, and perturbative stability coexist within a single controlled framework. We summarize these results in in Table 4.
VI Quantum complexity study of $H^{\mathrm{Tot}}$}
@@
Figure 13: Quantum complexity for the Hamiltonian $H^{\mathrm{Tot}}$, as defined in Eq. (1). (a) The behavior of entanglement for $L=18$ we obtain here remains notably distinct from that obtained for other KCMs discussed previously. Notably, in this case, in addition to the primary arch-like behavior, the presence of low-entangled states creates a sub-arch structure or an entanglement band. Similar patterns are observed in the plots shown in (b) and (c), where we plot the behavior of the Inverse Participation Ratio (IPR) for $L=18$ and the Stabilizer Rényi Entropy, $\mathcal{M}$ for $L=12$, respectively. In case of (b), the y-axis is rescaled by a factor $10^{-3}$. Finally, (d) illustrates the chaotic behavior of the same model for $L=20$.
@@
We devote this section to discuss the quantum complexity properties of the Hamiltonian $H_{\mathrm{Tot}}$, separately. Interestingly, in this case, we observe that although the model does not exhibit any fragmentation, its quantum complexity behavior shows a distinct feature from all the other models we have discussed so far. The first distinct feature we highlight here is the EE behavior with energy as shown in Fig. 13(a), displaying sub-arch-like splitting of the primary arch shape, resulting in an entanglement band-like structure. This behavior we find counterintuitive, and it does not translate immediately from all other observed behavior. To substantiate our claim, in Fig. 13(b), we also plot the Inverse Participation Ratio (IPR), defined as
@@
with $\{|i\rangle\}$ being the basis states in the symmetry reduced sector. The figure shows that IPR exhibits the same behavior as EE. The same imprint can be found in SRE as depicted in Fig. 13(c). Hence, although the system does not exhibit conventional fragmentation, it indicates that EE and other measures are not smooth functions of energy. Even in the bulk, there exist eigenstates that cause dips in the entanglement and create this band-like structure.
This observation motivates a closer investigation of the origin of these entanglement ripples. Notably, the total Hamiltonian can be written as
$H^{\mathrm{Tot}}=\sum_{i}\sigma^{x}_{i}-H^{\mathcal{N}(4)}.$ (13)
To systematically explore how these ripples emerge, we introduce a tunable Hamiltonian
$H^{\mathrm{Tot}}(\Delta)=\sum_{i}\sigma^{x}_{i}-\Delta\cdot H^{\mathcal{N}(4)}.$ (14)
Figure 14 shows the behavior of the entanglement entropy for different values of $\Delta$. For $\Delta=0$, the system reduces to the undressed model $H_{x}=\sum_{i}\sigma^{x}_{i}$”, which is integrable and exhibits localized entanglement behavior with well-separated bands. As soon as $\Delta$ is turned on, these localized bands begin to hybridize, leading to partial mixing between neighboring eigenstates. With increasing $\Delta$, the entanglement develops a ripple-like structure that we observe in the full $H^{\mathrm{Tot}}$ model at ($\Delta=1$) as also depcited in Fig. 13(a).
Interestingly, for sufficiently large values of $\Delta$, the ripples are completely washed out and the entanglement entropy becomes a smooth function of energy, consistent with ergodic behavior. This crossover highlights how weak integrability-breaking perturbations can generate nontrivial structure in eigenstate entanglement, even in the absence of strict fragmentation.
Figure 14: The entanglement entropy behavior of $H^{\mathrm{Tot}}(\Delta)$. This plot supports the behavior of EE obtained in Fig. 13. Here we analyze the entanglement entropy as a function of energy for the tunable Hamiltonian ($H^{\mathrm{Tot}}(\Delta)=\sum_{i}\sigma^{x}_{i}-\Delta H^{\mathcal{N}(4)})$. Starting from $\Delta=0$, corresponding to the undressed and integrable $\sigma_{i}^{x}$ model, the spectrum exhibits sharply defined entanglement bands. As $\Delta$ is increased, these bands begin to hybridize, leading to the emergence of ripple-like structures in the entanglement profile. For sufficiently large $\Delta$, the ripples are completely washed out and the entanglement entropy becomes a smooth function of energy.
@@ VII Discussion
In this work, we have investigated quantum complexity in a family of rule-based kinetically constrained models, examining both the challenges of state preparation and the growth of complexity under dynamical evolution. The family includes the celebrated quantum Game of Life and their variations. A key characteristic of this class of models is that it goes beyond the strict blockade condition encountered in the PXP family, which is realized through symmetric projectors. Instead, the quantum Hamiltonians considered here involve terms constructed from asymmetric projectors as well. We have shown how entanglement, nonstabilizerness, and quantum chaos interplay with each other in these models. In particular, we have unveiled the chaotic behavior through systematic symmetry resolutions of the models. However, there are cases where mere symmetry resolution has not revealed the chaotic behavior, as the Hilbert space exhibits dynamically disconnected sectors, which has resulted in Hilbert space fragmentation. Along with that, for some of the models, the fragmented subspaces have further hosted quantum many-body scars. This has given us a scope to extract non-trivial physics that exhibits quantum chaos, quantum many-body scars, as well as Hilbert space fragmentation out of a single family of quantum many-body models.
We have further characterized the fragmented subspaces by their resource generation capabilities. In particular, we have studied how the dimension of the Krylov subspaces correlates with the maximum entanglement entropy generation as well as their classical simulability aspects as quantified by Stabilizer Rényi entropy. Our analysis reveals that the static and dynamical behaviors of these models exhibit distinct capacities for quantum resource generation, which can be instrumental in their characterization. Hence, our study sheds light on the interplay between dynamical constraints, accessible Krylov subspaces, and the resulting complexity of quantum state preparation and evolution for structurally similar kinetically constrained models that can be instrumental for further fundamental as well as application-based studies. For example, in our work on QGL and related models, we have not found evidence of quantum many-body scars, which we plan to pursue more systematically in future work. Along with that, whether the quantum many-body scar or the fragmented phase can be applied for practical applications such as quantum sensing is another direction we also wish to explore. Another aim will be to characterize the fragment subspaces in terms of their Krylov Complexity behavior [66, 67, 68, 69]. Finally, the entanglement localization feature imprinted in all the quantum properties we obtained for the model $H^{\mathrm{Tot}}$ is another aspect that we will study in detail in our future work.